# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### v4

@@ 1 Introduction
Noise, traditionally viewed as an obstacle in machine learning applications Thulasidasan et al. (2019), is universal due to various factors such as environmental conditions, equipment calibration, and human activities Ormiston et al. (2020). In computer vision, noise can emerge at multiple stages. During image acquisition, for instance, imaging systems may introduce noise due to factors such as detector limitations, or electronic interference. This could manifest as electronic or thermal noise, leading to random variations in pixel values or color discrepancies in the captured images Sijbers et al. (1996). In specific imaging contexts, e.g. fluoroscopic imaging, additional noise components may be present, including photon scattering artifacts, uncorrelated quantum noise originating from photon statistical variations, and correlated noise associated with readout processes Zhang et al. (2014). Additionally, noise can also be introduced during the image preprocessing phase. Operations such as image resizing, filtering, interpolation or color space conversion are potential sources of noise Al-Shaykh & Mersereau (1998). For example, resizing might result in aliasing artifacts, non-linear interpolation methods, such as B-splines Hou & Andrews (1978), may lead to local morphological distortions in the image, and low-pass filtering could cause a loss of image detail and texture.
Prevailing literature typically assumes that noise adversely affects the task Sethna et al. (2001); Owotogbe et al. (2019). However, is this assumption always applicable? Our work seeks to thoroughly examine this critical question. We recognize that the vague definition of noise contributes to the uncertainty in identifying and characterizing it. One effective way to categorize different noises is through analysis of task complexity change  (Li, 2022). Leveraging the concept of task complexity, we can categorize noise into two types: positive noise (PN) and harmful noise (HN). PN reduces task complexity, whereas HN increases it, consistent with traditional views of noise.
Our work, which combines a theoretical analysis based on information theory with extensive empirical evaluation, reveals a surprising fact: the simple injection of noise into deep neural networks, when done in a principled manner, can significantly enhance model performance.
@@ 1.1 Scope and Contribution
Our work aims to investigate how various types of noise affect deep learning models. Specifically, the study focuses on three common types of noise, i.e., Gaussian noise, linear transform noise, and salt-and-pepper noise. Gaussian noise refers to random fluctuations that follow a Gaussian distribution in pixel values at the image level or latent representations. Linear transforms, on the other hand, refer to affine elementary transformations to the dataset of original images or latent representations. Salt-and-pepper noise is a kind of image distortion that adds random black or white values at the image level or to the latent representations.
This paper analyzes the impact of these types of noise on the performance of deep learning models for image classification tasks, including image classification and unsupervised domain adaptation. Two popular model families, Vision Transformers (ViTs) and Convolutional Neural Networks (CNNs), are considered in the study. Image classification, a foundational task in computer vision, involves predicting the class label of a given input image. In contrast, domain adaptation addresses scenarios where the training and test data originate from distinct distributions, commonly referred to as different domains. This comprehensive analysis aims to shed light on how noise impacts these critical tasks across different model architectures.
We start by presenting a comprehensive theoretical analysis of how these three types of noise impact deep models. Building on the theoretical foundation, we propose NoisyNN, a novel method designed to enhance the deep neural network performance on Image Classification and Domain Adaptation. To validate the effectiveness of NoisyNN, we conduct extensive experiments with ViTs and CNNs. Our empirical findings demonstrate the huge benefits of leveraging positive noise.
@@
  • We re-examined the conventional view that noise, by default, has a negative impact on deep learning models. Our theoretical analysis and experimental results show that noise can be a positive support for deep learning models on image classification tasks.
  • We introduce NoisyNN, an innovative approach that strategically utilizes positive noise. NoisyNN achieves state-of-the-art on various image classification and domain adaptation tasks.
  • Our study, along with the success of NoisyNN, prompts a re-visit of the role of noise in machine learning. This opens new avenues for future research to enhance deep learning model performance on various learning tasks.
@@ 1.2 Related Work
Positive Noise. Common knowledge often assumes noise to be detrimental to tasks. However, empirical evidence also suggests that some forms of noise could be beneficial. For instance, within the signal processing community, it has been shown that random noise can facilitate stochastic resonance, enhancing the detection of weak signals Benzi et al. (1981). In neuroscience, noise has been recognized for its potential to boost brain functionality McClintock (2002); Mori & Kai (2002).
Recent work by  Li (2022) marks a significant advance in the theoretical understanding of the characteristics of different noises. By employing information theory, they differentiate between beneficial “positive noise” and detrimental “pure noise”, based on their impact on task complexity. However, their analysis has three notable limitations: 1. it is confined to only the image space; 2. all experiments are conducted only on shallow models, far from the current best practices 3. it does not answer the practical question: how to create and leverage positive noise?
Our study aims to address these limitations. We make a significant extension to the positive-noise framework proposed by Li (2022). Our work not only confirms the presence of positive and harmful noise in embedding space but more importantly demonstrates that leveraging positive noise in deeper layers of the embedding space is often more effective (see Fig 2 b&d). Furthermore, we propose a practical approach to leverage the positive noise in deep models, we term it “NoisyNN”. This approach promises to unlock new potentials in the application of noise for enhancing neural network performance.
Data Augmentation. Data augmentation has been playing an important role in training deep vision models, as evidenced by its demonstrated effectiveness (Yang et al., 2023b). The general idea of data augmentation is to compose transformation operations that can be applied to the original data $x$ to create transformed data $x^{\prime}$ without severely altering the semantics. Common data augmentation range from simple techniques like random flip and crop (Krizhevsky et al., 2012) to more complex techniques like MixUp (Zhang et al., 2017), CutOut (DeVries & Taylor, 2017), AutoAugment (Cubuk et al., 2019), AugMix (Hendrycks et al., 2019) RandAugment (Cubuk et al., 2020) among others. More comprehensive reviews can be found in (Mumuni & Mumuni, 2022). The proposed approach is related to the research on data augmentation but stands apart due to its theoretical foundation. Our framework provides a more controlled and principled way to augment data, setting it apart from conventional methods, which often require substantial domain knowledge and ad-hoc design, as noted in (Cubuk et al., 2020).
Comparison with Manifold MixUp. Our NoisyNN shares some similarities with Manifold MixUp (Verma et al., 2019), a regularization technique designed for supervised image classification that extends the MixUp strategy to the embedding space by linearly interpolating embedding vectors $z_{i}$ (instead of images $x_{i}$) along with their corresponding labels $y_{i}$. However, there are several key differences. Unlike Manifold MixUp, which aims to flatten class representations through training on interpolated synthetic samples, our NoisyNN is grounded in a theoretical analysis of how noise injection impacts task complexity in view of information entropy, as introduced by (Li, 2022).
Additionally, we derived the optimal form of noise injection (Eq.20) within the linear transform noise design space, which Manifold MixUp does not provide. Procedurally, Manifold MixUp interpolates both embeddings and labels to generate synthetic samples, followed by training on these samples, as its theoretical foundation relies on modifying both features and labels. In contrast, our method perturbs only the embeddings and leaves the labels unchanged, as our theoretical analysis is based on un-interpolated labels. Investigating whether label interpolation could be integrated into our theoretical framework may be a promising avenue for future research.
@@ 2 Preliminary
In information theory, the entropy Shannon (2001) of a random variable $x$ is defined as:
[…]
where $p(x)$ is the distribution of the given variable $x$. The mutual information of two random discrete variables $(x,y)$ is denoted as Cover (1999):
@@
where $D_{KL}$ is the Kullback-Leibler divergence Kullback & Leibler (1951), and $p(x,y)$ is the joint distribution. The conditional entropy is defined as:
@@
The above definitions can be readily expanded to encompass continuous variables through the substitution of the sum operator with the integral symbol. In this work, the noise is denoted by $\bm{\epsilon}$ if without any specific statement.
@@
Since the entropy of task $\mathcal{T}$ is formulated, it is not difficult to define the entropy change when additional noise $\bm{\epsilon}$ is present Li (2022),
$\bigtriangleup S(\mathcal{T},\bm{\epsilon})=H(\mathcal{T})-H(\mathcal{T}|\bm{\epsilon})$ (4)
Formally, if the noise can help reduce the complexity of the task, i.e., $H(\mathcal{T}|\bm{\epsilon})<H(\mathcal{T})$ then the noise has positive support. Therefore, a noise $\bm{\epsilon}$ is defined as positive noise (PN) when the noise satisfies $\bigtriangleup S(\mathcal{T},\bm{\epsilon})>0$. On the contrary, when $\bigtriangleup S(\mathcal{T},\bm{\epsilon})\leq 0$, the noise is considered as the conventional noise and named harmful noise (HN).
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\bm{\epsilon})>0&\bm{\epsilon}\ \mathrm{is\ positive}\ \mathrm{noise}\\ \bigtriangleup S(\mathcal{T},\bm{\epsilon})\leq 0&\bm{\epsilon}\ \mathrm{is\ harmful}\ \mathrm{noise}\end{matrix}\right.$ (5)
Moderate Model Assumption: Positive noise may not be effective for deep models facing severe issues, such as overfitting, where the model starts to memorize random fluctuations in the data rather than learning the underlying patterns. In such cases, positive noise is unlikely to provide meaningful support in improving the model’s performance.
Figure 1: An overview of the proposed method: Above the black line is the standard pipeline for image classification, where the deep model can be either a CNN or a ViT. Noise is injected into a randomly selected layer of the model, as indicated by the blue arrow. Once the layer is selected, it remains fixed during both training and inference.
@@ 3 Methods
The idea of exploring the influence of noise on deep models is straightforward: we add noise to the latent representations. The framework is depicted in Fig. 1. This is a universal framework for classification tasks where there are different options for deep models, such as CNNs and ViTs. Through the simple operation of injecting noise into a randomly selected layer, a model has the potential to gain additional information to reduce task complexity, thereby improving its performance. It is sufficient to inject noise into a single layer instead of multiple layers since it imposes a regularization on other layers of the model simultaneously.
@@
For a classification problem, the dataset $(\bm{X},\bm{Y})$ can be regarded as samplings derived from $\mathit{D}_{\mathcal{X},\mathcal{Y}}$, where $\mathit{D}_{\mathcal{X},\mathcal{Y}}$ is some unknown joint distribution of data points and labels from feasible space $\mathcal{X}$ and $\mathcal{Y}$, i.e., $(\bm{X},\bm{Y})\sim\mathit{D}_{\mathcal{X},\mathcal{Y}}$ Shalev-Shwartz & Ben-David (2014). Hence, given a set of $k$ data points $\bm{X}=\left\{X_{1},X_{2},...,X_{k}\right\}$, the label set $\bm{Y}=\left\{Y_{1},Y_{2},...,Y_{k}\right\}$ is regarded as sampling from $\bm{Y}\sim\mathit{D}_{\mathcal{Y}|\mathcal{X}}$. The complexity of $\mathcal{T}$ on dataset $\bm{X}$ is formulated as:
$H(\mathcal{T};\bm{X})=H(\bm{Y};\bm{X})-H(\bm{X})$ (6)
Accordingly, the operation of adding noise at the image level can be formulated as Li (2022):
@@
$\left\{\begin{matrix}H(\mathcal{T};\bm{X}+\bm{\epsilon})=-\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Y}|\bm{X}+\bm{\epsilon})\log{p(\bm{Y}|\bm{X}+\bm{\epsilon})}&\bm{\epsilon}\ \mathrm{is\ additive\ noise}\\ H(\mathcal{T};\bm{X}\bm{\epsilon})=-\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Y}|\bm{X}\bm{\epsilon})\log{p(\bm{Y}|\bm{X}\bm{\epsilon})}&\bm{\epsilon}\ \mathrm{is\ multiplicative\ noise}\end{matrix}\right.$ (7)
@@
Inspired by the previous work Li (2022), given a set of $k$ image embeddings $\bm{Z}=\left\{Z_{1},Z_{2},...,Z_{k}\right\}$, the label set $\bm{Y}=\left\{Y_{1},Y_{2},...,Y_{k}\right\}$ is regarded as sampling from $\bm{Y}\sim\mathit{D}_{\mathcal{Y}|\mathcal{Z}}$. The complexity of $\mathcal{T}$ on dataset $\bm{Z}$ is formulated as:
$H(\mathcal{T};\bm{Z})\coloneqq H(\bm{Y};\bm{Z})-H(\bm{Z})$ (8)
The operation of proactively injecting noise in the latent space is defined as:
@@
$\left\{\begin{matrix}H(\mathcal{T};\bm{Z}+\bm{\epsilon})\coloneqq H(\bm{Y};\bm{Z}+\bm{\epsilon})-H(\bm{Z})&\bm{\epsilon}\ \mathrm{is\ additive\ noise}\\ H(\mathcal{T};\bm{Z}\bm{\epsilon})\coloneqq H(\bm{Y};\bm{Z}\bm{\epsilon})-H(\bm{Z})&\bm{\epsilon}\ \mathrm{is\ multiplicative\ noise}\end{matrix}\right.$ (9)
@@
where $\bm{Z}$ are the embeddings of the images. The definition of Eq. 9 differs from the conventional definition, as our method injects the noise into the latent representations instead of the original images. The Gaussian noise is additive, the linear transform noise is also additive, while the salt-and-pepper is multiplicative.
@@
$\begin{matrix}H(\mathcal{T};\bm{Z}+\bm{\epsilon})=H(\bm{Y};\bm{Z}+\bm{\epsilon})-H(\bm{Z})\end{matrix}.$ (10)
We assume that both $\bm{Z}$ and $\bm{Y}$ follow a multivariate normal distribution. Additionally, we can transform the distributions of $\bm{Z}$ and $\bm{Y}$ to make them (approximately) follow the multivariate normal distribution, even if they initially do not Box & Cox (1964) Feng et al. (2014). According to the definition in Equation 4, the entropy change with Gaussian noise is:
$\begin{split}\bigtriangleup S(\mathcal{T},\bm{\epsilon})=&H(\bm{Y};\bm{Z})-H(\bm{Z})-(H(\bm{Y};\bm{Z}+\bm{\epsilon})-H(\bm{Z}))\\ =&H(\bm{Y};\bm{Z})-H(\bm{Y};\bm{Z}+\bm{\epsilon})\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Z}}||\bm{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|}{|\bm{\Sigma_{Z+\epsilon}}||\bm{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{ZY}}|}\\ =&\frac{1}{2}\log\frac{1}{(1+\sigma^{2}_{\epsilon}{\textstyle\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}})(1+\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))}})}\end{split}$ (11)
where $\lambda=\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}$, $\sigma^{2}_{\epsilon}$ is the variance of the Gaussian noise, $\mathrm{cov}(Z_{i},Y_{i})$ is the covariance of sample pair ${X_{i},Y_{i}}$, $\sigma^{2}_{Z_{i}}$ and $\sigma^{2}_{Y_{i}}$ are the variance of data sample $Z_{i}$ and data label $Y_{i}$, respectively. The detailed derivations can be found in section 1.1.2 of the supplementary. Given a dataset, the variance of the Gaussian noise, and statistical properties of data samples and labels control the entropy change, we define the function:
@@
Since ${\epsilon}^{2}\geq 0$ and $\lambda\geq 0$, $\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i})=\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}(1-\rho_{Z_{i}Y_{i}}^{2})\geq 0$, where $\rho_{Z_{i}Y_{i}}$ is the correlation coefficient, the sign of $M$ is negative. Consequently, we conclude that the injection of Gaussian noise into the embedding space is harmful to the task. Detailed derivations can be found in App sec. B.
Linear Transform Noise. This type of noise is obtained by elementary transformation of the features matrix, i.e., $\bm{\epsilon}=Q\bm{X}$, where $Q$ is a linear transformation matrix. We name the $Q$ the quality matrix since it controls the property of linear transform noise and determines whether positive or harmful. In the linear transform noise injection in the latent space case, the complexity of the task is:
$\begin{split}H(\mathcal{T};\bm{Z}+Q\bm{Z})=H(\bm{Y};\bm{Z}+Q\bm{Z})-H(\bm{Z})\end{split}$ (13)

### v5

@@ 1 Introduction
Noise, conventionally regarded as a hurdle in machine learning and deep learning tasks, is universal and unavoidable due to various reasons, e.g., environmental factors, instrumental calibration, and human activities [40] [58]. In computer vision, noise can be generated from different phases: (1) Image Acquisition: Noise can arise from a camera sensor or other imaging device [53]. For example, electronic or thermal noise in the camera sensor can result in random pixel values or color variations that can be visible in the captured image. (2) Image Preprocessing: Noise can be introduced during preprocessing steps such as image resizing, filtering, or color space conversion [1]. For example, resizing an image can introduce aliasing artifacts, while filtering an image can result in the loss of detail and texture. (3) Feature Extraction: Feature extraction algorithms can be sensitive to noise in the input image, which can result in inaccurate or inconsistent feature representations [2]. For example, edge detection algorithms can be affected by noise in the image, resulting in false positives or negatives. (4) Algorithms: algorithms used for computer vision tasks, such as object detection or image segmentation, can also be sensitive to noise in the input data [6]. Noise can cause the algorithm to learn incorrect patterns or features, leading to poor performance.
Since noise is an unavoidable reality in engineering tasks, existing works usually make the assumption that noise has a consistently negative impact on the current task [48] [41]. Nevertheless, is the above assumption always valid? As such, it is crucial to address the question of whether noise can ever have a positive influence on deep learning models. This work aims to provide a comprehensive answer to this question, which is a pressing concern in the deep learning community. We recognize that the imprecise definition of noise is a critical factor leading to the uncertainties surrounding the identification and characterization of noise. To address these uncertainties, an in-depth analysis of the task’s complexity is imperative for arriving at a rigorous answer. By using the definition of task complexity, it is possible to categorize noise into two distinct categories: positive noise (PN) and harmful noise (HN). PN decreases the complexity of the task, while HN increases it, aligning with the conventional understanding of noise.
@@ 1.1 Scope and Contribution
Our work aims to investigate how various types of noise affect deep learning models. Specifically, the study focuses on three common types of noise, i.e., Gaussian noise, linear transform noise, and salt-and-pepper noise. Gaussian noise refers to random fluctuations that follow a Gaussian distribution in pixel values at the image level or latent representations in latent space [47]. Linear transforms, on the other hand, refer to affine elementary transformations to the dataset of original images or latent representations, where the transformation matrix is row equivalent to an identity matrix [57]. Salt-and-pepper noise is a kind of image distortion that adds random black or white values at the image level or to the latent representations [7].
This paper analyzes the impact of these types of noise on the performance of deep learning models for image classification and domain adaptation tasks. Two popular model families, Vision Transformers (ViTs) and Convolutional Neural Networks (CNNs), are considered in the study. Image classification is one of the most fundamental tasks in computer vision, where the goal is to predict the class label of an input image. Domain adaptation is a practically meaningful task where the training and test data come from different distributions, also known as different domains. By investigating the effects of different types of noise on ViTs and CNNs for typical deep learning tasks, the paper provides insights into the influences of noises on deep models. The findings presented in this paper hold practical significance for enhancing the performance of various types of deep learning models in real-world scenarios.
@@
  • We re-examined the conventional view that noise, by default, has a negative impact on deep learning models. Our theoretical analysis and experimental results show that noise can be a positive support for deep learning models and tasks.
  • We implemented extensive experiments with different deep models, such as CNNs and ViTs, and on different deep learning tasks. Empowered by positive noise, we achieved state-of-the-art (SOTA) results in the experiments presented in this paper.
  • Instead of operating on the image level, our injecting noise operations are performed in the latent space. We theoretically analyze the difference between injecting noise on the image level and in the latent space.
  • The theory and framework of reducing task complexity via positive noise in this work can be applied to any deep learning architecture. There is great potential for exploring the application of positive noise in other deep-learning tasks beyond the image classification and domain adaptation tasks examined in this study.
@@ 1.2 Related Work
Positive Noise In fact, within the signal-processing society, it has been demonstrated that random noise helps stochastic resonance improve the detection of weak signals [4]. Noises can have positive support and contribute to less mean square error compared to the best linear unbiased estimator when the mixing probability distribution is not in the extreme region [45]. Also, it has been reported that noise could increase the model generalization in natural language processing (NLP) [44]. Recently, the perturbation, a special case of positive noise, has been effectively utilized to implement self-refinement in domain adaptation and achieved state-of-the-art performance [57]. The latest research shows that by proactively adding specific noise to partial datasets, various tasks can benefit from the positive noise [30]. Besides, noises are found to be able to boost brain power and be useful in many neuroscience studies [37] [39].
Deep Model Convolutional Neural Networks have been widely used for image classification, object detection, and segmentation tasks, and have achieved impressive results [28][22]. However, these networks have limitations in terms of their ability to capture long-range dependencies and extract global features from images. Recently, Vision Transformer has been proposed as an alternative to CNNs [15]. ViT relies on self-attention mechanisms and a transformer-based architecture to enable global feature extraction and modeling of long-range dependencies in images [61]. The attention mechanism allows the model to focus on the most informative features of the input image, while the transformer architecture facilitates information exchange between different parts of the image. ViT has demonstrated impressive performance on a range of image classification tasks and has the potential to outperform traditional CNN-based approaches. However, ViT currently requires a large number of parameters and training data to achieve state-of-the-art results, making it challenging to implement in certain settings [68].
@@ 2 Preliminary
In information theory, the entropy [50] of a random variable $x$ is defined as:
$H(x)=\left\{\begin{matrix}-\int p(x)\log p(x)dx&\quad\ \mathrm{if}\ x\ \mathrm{is\ continuous}\\ -\sum_{x}p(x)\log p(x)&\mathrm{if}\ x\ \mathrm{is\ discrete}\end{matrix}\right$. (1)
where $p(x)$ is the distribution of the given variable $x$. The mutual information of two random discrete variables $(x,y)$ is denoted as [9]:
@@
where $D_{KL}$ is the Kullback–Leibler divergence [26], and $p(x,y)$ is the joint distribution. The conditional entropy is defined as:
@@
The above definitions can be readily expanded to encompass continuous variables through the substitution of the sum operator with the integral symbol. In this work, the noise is denoted by $\boldsymbol{\epsilon}$ if without any specific statement.
@@
Since the entropy of task $\mathcal{T}$ is formulated, it is not difficult to define the entropy change when additional noise $\boldsymbol{\epsilon}$ is present [30],
$\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=H(\mathcal{T})-H(\mathcal{T}|\boldsymbol{\epsilon})$ (4)
Formally, if the noise can help reduce the complexity of the task, i.e., $H(\mathcal{T}|\boldsymbol{\epsilon})<H(\mathcal{T})$, then the noise has positive support. Therefore, a noise $\boldsymbol{\epsilon}$ is defined as positive noise (PN) when the noise satisfies $\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0$. On the contrary, when $\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0$, the noise is considered as the conventional noise and named harmful noise (HN).
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0&\boldsymbol{\epsilon}\ \mathrm{is\ positive}\ \mathrm{noise}\\ \bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0&\boldsymbol{\epsilon}\ \mathrm{is\ harmful}\ \mathrm{noise}\end{matrix}\right$. (5)
Moderate Model Assumption: The positive noise may not work for deep models with severe problems. For example, the model is severely overfitting where models begin to memorize the random fluctuations in the data instead of learning the underlying patterns. In that case, the presence of positive noise will not have a positive impact on improving the models’ performance.
Figure 1: An overview of the proposed method. Above the black line is the standard pipeline for image classification. The deep model can be CNNs or ViTs. The noise is injected into a randomly chosen layer of the model represented by the blue arrow.
@@ 3 Methods
The idea of exploring the influence of noise on the deep models is straightforward. The framework is depicted in Fig. 1. This is a universal framework where there are different options for deep models, such as CNNs and ViTs. Through the simple operation of injecting noise into a randomly selected layer, a model has the potential to gain additional information to reduce task complexity, thereby improving its performance. It is sufficient to inject noise into a single layer instead of multiple layers since it imposes a regularization on multiple layers simultaneously.
@@
For a classification problem, the dataset $(\boldsymbol{X},\boldsymbol{Y})$ can be regarded as samplings derived from $\mathit{D}_{\mathcal{X},\mathcal{Y}}$, where $\mathit{D}_{\mathcal{X},\mathcal{Y}}$ is some unknown joint distribution of data points and labels from feasible space $\mathcal{X}$ and $\mathcal{Y}$, i.e., $(\boldsymbol{X},\boldsymbol{Y})\sim\mathit{D}_{\mathcal{X},\mathcal{Y}}$ [49]. Hence, given a set of $k$ data points $\boldsymbol{X}=\left\{X_{1},X_{2},...,X_{k}\right\}$, the label set $\boldsymbol{Y}=\left\{Y_{1},Y_{2},...,Y_{k}\right\}$ is regarded as sampling from $\boldsymbol{Y}\sim\mathit{D}_{\mathcal{Y}|\mathcal{X}}$. The complexity of $\mathcal{T}$ on dataset $\boldsymbol{X}$ is formulated as:
$H_{\mathcal{T}}(\boldsymbol{X})\coloneqq H(\boldsymbol{Y}\mid\boldsymbol{X})$ (6)
Accordingly, the operation of adding noise at the image level can be formulated as [30]:
@@
$\left\{\begin{matrix}H_{\mathcal{T}}(\boldsymbol{X}+\boldsymbol{\epsilon})=-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{X}+\boldsymbol{\epsilon})\log{p(\boldsymbol{Y}|\boldsymbol{X}+\boldsymbol{\epsilon})}&\boldsymbol{\epsilon}\ \mathrm{is\ additive\ noise}\\ H_{\mathcal{T}}(\boldsymbol{X}\boldsymbol{\epsilon})=-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{X}\boldsymbol{\epsilon})\log{p(\boldsymbol{Y}|\boldsymbol{X}\boldsymbol{\epsilon})}&\boldsymbol{\epsilon}\ \mathrm{is\ multiplicative\ noise}\end{matrix}\right$. (7)
@@
Inspired by the previous work [30], given a set of $k$ image embeddings $\boldsymbol{Z}=\left\{Z_{1},Z_{2},...,Z_{k}\right\}$, the label set $\boldsymbol{Y}=\left\{Y_{1},Y_{2},...,Y_{k}\right\}$ is regarded as sampling from $\boldsymbol{Y}\sim\mathit{D}_{\mathcal{Y}|\mathcal{Z}}$. The complexity of $\mathcal{T}$ on dataset $\boldsymbol{Z}$ is formulated as:
$H_{\mathcal{T}}(\boldsymbol{Z})\coloneqq H(\boldsymbol{Y},\boldsymbol{Z})-H(\boldsymbol{Z})$ (8)
We emphasize that $H(\boldsymbol{Z})$ refers to the joint entropy of all embeddings in the dataset. Our formulation captures the dataset-level information content rather than per-sample entropy. In this work, we freeze the vision encoder, so the marginal embedding entropy $H(\boldsymbol{Z})$ is kept as a fixed baseline. We then define the task complexity under proactive noise injection in the latent space as:
@@
$\left\{\begin{matrix}H_{\mathcal{T}}(\boldsymbol{Z}+\boldsymbol{\epsilon})\coloneqq H(\boldsymbol{Y},\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z})&\boldsymbol{\epsilon}\ \mathrm{is\ additive\ noise}\\ H_{\mathcal{T}}(\boldsymbol{Z}\boldsymbol{\epsilon})\coloneqq H(\boldsymbol{Y},\boldsymbol{Z}\boldsymbol{\epsilon})-H(\boldsymbol{Z})&\boldsymbol{\epsilon}\ \mathrm{is\ multiplicative\ noise}\end{matrix}\right$. (9)
@@
where $\boldsymbol{Z}$ are the embeddings of the images. The definition of Eq. 9 differs from the conventional definition, as our method injects the noise into the latent representations instead of the original images. The Gaussian noise is additive, the linear transform noise is also additive, while the salt-and-pepper noise is multiplicative.
@@
$\begin{matrix}H_{\mathcal{T}}(\boldsymbol{Z}+\boldsymbol{\epsilon})=H(\boldsymbol{Y},\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z})\end{matrix}.$ (10)
We assume that both $\boldsymbol{Z}$ and $\boldsymbol{Y}$ follow a multivariate normal distribution. Additionally, we can transform the distributions of $\boldsymbol{Z}$ and $\boldsymbol{Y}$ to make them (approximately) follow the multivariate normal distribution, even if they initially do not [5] [17]. According to the definition in Equation 4, the entropy change with Gaussian noise is:
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H_{\mathcal{T}}(\boldsymbol{Z})-H_{\mathcal{T}}(\boldsymbol{Z}+\boldsymbol{\epsilon})\\ =&H(\boldsymbol{Y},\boldsymbol{Z})-H(\boldsymbol{Z})-(H(\boldsymbol{Y},\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z}))\\ =&H(\boldsymbol{Y},\boldsymbol{Z})-H(\boldsymbol{Y},\boldsymbol{Z}+\boldsymbol{\epsilon})\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{ZY}}|}\\ =&\frac{1}{2}\log\frac{1}{(1+\sigma^{2}_{\epsilon}{\textstyle\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}})(1+\lambda{\sum_{i=1}^{k}\frac{\mathrm{cov}^{2}(Z_{i},Y_{i})}{\sigma^{2}_{X_{i}}(\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i}))}})}\end{split}$ (11)
where $\lambda=\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}$, $\sigma^{2}_{\epsilon}$ is the variance of the Gaussian noise, $\mathrm{cov}(Z_{i},Y_{i})$ is the covariance of sample pair $X_{i},Y_{i}$}, $\sigma^{2}_{Z_{i}}$ and $\sigma^{2}_{Y_{i}}$ are the variance of data sample $Z_{i}$ and data label $Y_{i}$, respectively.
The detailed derivations can be found in section 1.1.2 of the supplementary. Given a dataset, the variance of the Gaussian noise, and statistical properties of data samples and labels control the entropy change, we define the function:
@@
Since ${\epsilon}^{2}\geq 0$ and $\lambda\geq 0$, $\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}-\mathrm{cov}^{2}(Z_{i},Y_{i})=\sigma^{2}_{Z_{i}}\sigma^{2}_{Y_{i}}(1-\rho_{Z_{i}Y_{i}}^{2})\geq 0$, where $\rho_{Z_{i}Y_{i}}$ is the correlation coefficient, the sign of $M$ is negative. We can conclude that Gaussian noise injected into the latent space is harmful to the task. More details and the Gaussian noise added to the image level are provided in the supplementary.
Linear Transform Noise This type of noise is obtained by elementary transformation of the features matrix, i.e., $\boldsymbol{\epsilon}=Q\boldsymbol{X}$, where $Q$ is a linear transformation matrix. We name the $Q$ the quality matrix since it controls the property of linear transform noise and determines whether positive or harmful. In the linear transform noise injection in the latent space case, the complexity of the task is:
$\begin{split}H_{\mathcal{T}}(\boldsymbol{Z}+Q\boldsymbol{Z})=H(\boldsymbol{Y},\boldsymbol{Z}+Q\boldsymbol{Z})-H(\boldsymbol{Z})\end{split}$ (13)